### v1

@@ Abstract
Context: The POSiCS camera is a handheld, small field-of-view gamma camera developed for multipurpose use in radio-guided surgery (RGS), with sentinel lymph node biopsy (SLNB) as its benchmark application. This compact and lightweight detector (weighing approximately ) can map tissues labeled with nanocolloids and guide surgeons to the location of target lesions. By enabling intraoperative visualization in close proximity to the surgical field, its primary objective is to minimize surgical interventional invasiveness and operative time, thereby enhancing localization accuracy and reducing the incidence of post-operative complications. The design and components of the POSiCS camera emphasize ergonomic handling and compactness, providing, at the same time, rapid image formation and a spatial resolution of a few millimeters. These features are compatible with routine operating-room workflow, including wireless communication with the computer and a real-time display to support surgeon decision-making.
Methods: The camera was developed and assessed in two interchangeable hardware configurations: high-sensitivity and high-resolution, to provide flexibility according to application-specific requirements. Performance characterization employed Flood phantom (flat) sources of and point sources of . The response to was also investigated, given its increasing role in theranostics. Tests on the device were aimed at assessing the main performance parameters, such as extrinsic spatial resolution, energy resolution, and sensitivity. Moreover, the camera shielding was also evaluated, as well as the device’s response to gamma rays of different energy.
Results: The spatial resolution measured at a source–detector distance of was for the high sensitivity mode and for the high resolution mode. The system sensitivity at was evaluated as (high sensitivity) and (high resolution). For both working modes, we report an energy resolution of approximately 20%, even though the high-resolution collimator exhibits an increased scattered component due to the larger amount of tungsten.
@@
Intraoperative Gamma Camera, Nuclear Medicine, Radio Guided Surgery, Sentinel Lymph Nodes Biopsy, Radio Occult Lesion Localization, Silicon Photomultipliers1 Introduction
@@
Radio-guided surgery (RGS), including intraoperative techniques such as sentinel lymph node biopsy (SLNB) and radioguided occult lesion localization (ROLL), is a widely adopted approach, particularly in the surgical management of breast cancer and cutaneous melanoma. [RGS-breast-cancer, Mariani811]. Despite the availability of well-established alternatives to RGS (such as Fluorescence-Guided Surgery (FGS) or Ultrasound-Guided Surgery), the preoperative administration of radio-labeled nanocolloids remains one of the most widely used techniques for intraoperative localization of oncological lesions. This is due to the extensive range of gamma rays in living tissues, which allows for the direct detection and localization of activity hot-spots from outside the patient’s body [RGS-breast-cancer, Mariani811, Arabi_2024].
@@
The state-of-the-art practice in SLNB and ROLL is based on 1D probes to localize lesions or sentinel lymph nodes (SLNs) [Farnworth2023]. Although this approach has been proven reliable by clinically comparing preoperative lymphoscintigraphies with the number of SLNs localized through a probe-based search, probes still have some disadvantages. For this reason, an increasing number of 2D Intraoperative Gamma Cameras (IGCs), also known as Small Field of View (SFOV) gamma cameras, have been developed in the last decades [Stoffels2012, Massari2018]. This trend has been thoroughly studied in recent years, thanks to the comprehensive reviews in the field of IGCs, such as [Tsuchimochi_2013] and a more recent follow-up [Farnworth2023]. The main advantages of cameras are their imaging capabilities, which allow surgeons to localize hot spots near the ”blast zone” (radiotracer injection site) and provide a method to verify post-excision images to ensure that the target lymph node or lesion was entirely resected [Hellingman_2016, Paredes2008]. Additionally, imaging is intuitive and does not require the specific and lengthy training needed to become proficient with 1D probes. It has also been shown that 2D imaging of oncological lesions can reduce surgery duration, as it facilitates the localization of target tissues more quickly [Hellingman_2016, soluri2006].
The device is equipped with two interchangeable collimators: a short low-energy high-sensitivity (LEHS) collimator and a longer low-energy high-resolution (LEHR) collimator, enabling dual operating modes that can be selected intraoperatively to adapt to different clinical needs during surgery. Optimized for detecting photons at , corresponding to the emission energy of [KleinJan2013], the standard radionuclide used in RGS, the POSiCS camera exhibits enhanced performance relative to existing systems. This improvement is due in part to a novel position-sensitive silicon photomultiplier (SiPM) technology developed by Fondazione Bruno Kessler (FBK). The device performs center-of-gravity (CoG) light-position reconstruction, achieving a nominal spatial resolution of about [Acerbi2024, raiola2025spatialresolution]. Thanks to a “smart-channel” configuration, the SiPM array requires only readout channels over a area for a LG-SiPMs array [Acerbi2024]. This architecture reduces data throughput, provides fast timing, and enables wireless communication with the processing unit, thereby eliminating cumbersome cables during surgery. Moreover, the use of LG-SiPMs facilitates a battery-powered module, owing to the low power consumption of solid-state photodetectors [Acerbi2024, Gola2013ANA, Acerbi2023].
@@
In this article, we present the results of the first performance studies on the prototype POSiCS module. In Section 2, we describe the main components of the camera and its advantages in the field of RGS. In Section 3, we present the experimental methods adopted to measure the main performance parameters of the camera with both collimators. The tests were inspired by the 2023 NEMA standards for gamma camera qualification [NEMA2023]. However, as these standards are primarily designed for large clinical scanners, only a subset of the proposed measurements and procedures applies to our device, given its specific design and use case. In Section 4, the results of the tests are displayed and commented. Conclusively, in Section 4.6, we provide an overall discussion of the obtained results, while in Section 5, we offer a brief outlook and prospects.
2 The POSiCS camera
2.1 General operation
The POSiCS camera is a handheld gamma-ray imaging detector, based on a LYSO:Ce scintillator and a 33 array of position-sensitive SiPMs (LG-SiPMs).
For each detected event , the 8 signals from the SiPMs are digitized by an onboard ADC and processed by a microcontroller, which calculates the (x, y) position of the gamma-ray interaction. The position (x,y coordinates), the energy of the event (sum of the channels’ charge output), and the time stamp information are transmitted wirelessly to a PC, where image correction and quality enhancement algorithms are applied to produce the final image.
@@ 2.1 General operation
The current camera module has a total length of 147 mm (155 mm with the LEHR collimator mounted), with a Field of View (FOV) of . From this, we define a Useful Field of View (UFOV) of . The module total weight is 313 g (with the LEHS collimator) and 381 g (with the LEHR collimator). Further technical details about the camera are found in the POSiCS camera’s technical paper [Acerbi2025].
2.2 Collimator geometry
@@
The POSiCS camera is equipped with two interchangeable parallel-hole collimators: one targeting high spatial resolution (LEHR) for surgical procedures that require precise identification of target lesion margins, and the other designed to achieve high sensitivity (LEHS) for faster imaging times, even at low injected activities. The different parameters that qualify both objects were chosen through theoretical calculations and simulations using GATE [GATE2004]. Due to its high atomic number (implying a short attenuation length) and good resistance to shocks [Farnworth2023], tungsten was chosen as a material. All the collimator parameters were chosen to ensure a septal penetration probability of less than [collimator-design, cherry_ch14]. The collimator was designed with external shielding walls that extend beyond the collimator holes and support the scintillator, reducing the amount of gamma rays entering from the side of the camera.
2.3 LYSO scintillator approach
@@
The radiation-sensitive component of the POSiCS camera is a pixelated LYSO:Ce scintillator. This Lutetium-based crystal is typically used in Positron Emission Tomography (PET) scanners due to its fast decay time (), large light yield (), and high density [Kimble_2002]. The camera was therefore built using a very thin crystal (a few millimeters thick), keeping a high photoelectric absorption probability at . The scintillator is pixelated in a 3030 matrix. Each pixel is coated with a BaSO4 reflector to increase light collection on the light detector.
@@ 2.4 LG-SiPMs photodetector
In the current context of light detectors for IGCs, POSiCS proposes an innovative approach by moving towards PS-SiPMs, leveraging the new Linearly Graded SiPMs technology (LG-SiPMs) designed and produced by FBK [FBK-LG-SiPM, Acerbi2024, Acerbi2024_ps]. Through an array of 3x3 LG-SiPMs with peak sensitivity at , an active area of about per sensor, and microcell pitch of , it is possible to reconstruct the position of each scintillation event, through a weighted charge (and amplitude) distribution towards the output channels. Each Single Photon Avalanche Diode (SPAD) composing the SiPM is connected to a resistive network, which performs a light spot CoG reconstruction, encoded in the relative amount of charge output by each readout channel [Acerbi2025, Acerbi2024]. This Anger CoG reconstruction is typically used between SiPMs or PMTs in standard gamma cameras, but this approach is implemented at the single microcell level within the POSiCS camera. The 33 array configuration, with the ”smart channel” approach (see [Acerbi2024]), allows CoG reconstruction with only 8 readout channels: 4 encoding for the x coordinate () and 4 reconstructing the y coordinate (). More precisely, position reconstruction is performed as follows:
This weighted position calculation is performed directly within the camera digital electronics, which in turn sends the reconstructed positions through a wireless connection. This photon detector demonstrated exceptional intrinsic precision, achieving a resolution of a few hundred microns [Acerbi2024, raiola2025spatialresolution].
@@
2.5 Approach to image reconstruction
When a scintillation event occurs, the output charge from each of the eight channels is amplified by dedicated transimpedance amplifiers (TIAs). The summed signal is simultaneously processed by a comparator, which enables signal integration only if a user-defined threshold is exceeded. Once this condition is met, the charge of each channel is integrated and digitized by an eight-channel analog-to-digital converter (ADC). The microcontroller (C) then calculates the total event energy (in ADC counts), records the event timestamp, and computes the - and -coordinates using Equation 1. In addition, the C provides slow-control (SC) data to the system, including the bias voltage status, absorbed current, module temperature, battery status, and other relevant operational parameters. A block diagram of the analog and digital electronics of the POSiCS wireless module is shown in Figure 2. Each event consists of four parameters: the and scintillation light center-of-gravity (CoG) coordinates, the energy in analog-to-digital converter (ADC) counts, and the timestamp , represented as . The microcontroller (C) transmits these events to an external computer via an ESP32 module, in packets containing 4000 events. The user can retrieve these packets through HTTP GET requests, as well as access the slow-control (SC) information.111The camera settings are also transmitted through the same ESP32 module via HTTP POST requests, which configure the digital-to-analog converters (DACs). These commands include the bias voltage setting, comparator threshold level, and the integration delay and time width.
Reconstructed images are corrected using a three-step calibration and correction procedure. First, once the incident position of a gamma ray is determined (x,y CoG coordinates), the event is assigned to the corresponding camera pixel associated with the scintillator crystal encompassing the interaction position, whose physical dimensions are precisely known (linearity correction). This spatial mapping is established through a flood-field calibration scan, which serves as the primary calibration step before operating the camera.
Second, a pixel-by-pixel energy calibration is performed. In this step, a scaling coefficient is applied to convert the SiPMs charge response into a physical energy. Due to non-uniform light collection across certain regions of the FOV, as well as slight variations in the gain and response characteristics of the SiPMs forming the optical readout array, each pixel is assigned an individually determined conversion coefficient from ADC counts (digitized charge) to a physical energy in keV. Following this step, a user-defined energy window can be applied in software to selectively retain events consistent with photoelectric interactions, thereby suppressing contributions from Compton-scattered events that occur either within the LYSO crystal or in the surrounding environment.
@@ 2.5 Approach to image reconstruction
With this system, it is possible to live display images with a delay in the order of 1 or 2 seconds. Final images are usually enhanced with a Gaussian kernel, to smooth statistical variations in the image shape (especially useful in low-activity scenarios, where high statistics are difficult to achieve). The convolution kernel is defined as:
Following the 2023 guideline presented in the NEMA standards for the validation of gamma cameras [NEMA2023]. To conclude, the smoothed image is treated with a spline interpolation of order 2, with a zooming factor of 6. A scheme of the full data reconstruction and image enhancement pipeline is presented in Figure 2.
3 Experimental methods
@@
To assess spatial resolution, a point source, defined as having a diameter smaller than that of the scintillator pixel, was necessary. Due to the practical limitations of producing such a source with , a solid source was employed. This isotope emits at 122 keV, serving as a suitable surrogate for the 140.5 keV photons of .
@@
3.1 Extrinsic Spatial resolution
@@
The uncertainty on the resolution is the summation in quadrature of the statistical error (obtained through the covariance matrix of the fit) and a systematic uncertainty. Systematics is defined as the variation in resolution due to the different behavior of the camera in other areas of the FOV. The systematic uncertainty is evaluated through a scan of 15 positions on the device’s FOV with the Cobalt point source, acquiring 100’000 events at each position. Each reconstructed spot is fitted with a Gaussian profile, from which the FWHM is extracted. The standard deviation of this sample of 15 measured resolution values represents the systematic error of the resolution measurement.
The measurement at the center of the FOV is repeated for different source distances and under two conditions: with and without scattering material between the source and the camera head. The chosen scattering material is Polymethyl methacrylate (PMMA). This approach enables us to estimate the expected resolution in biological tissues, whose density is comparable to that of PMMA. The obtained resolution is also compared with the theoretical spatial resolution [AngerCamera, cherry_ch14, xcom2010]:
3.2 Extrinsic sensitivity
The camera’s sensitivity is defined as the registered trigger rate divided by the activity of the source within the camera’s FOV. This parameter is strongly dependent on the geometry of the collimator, the efficiency of the scintillator crystal (i.e., the probability of absorbing a gamma ray through photoelectric effect at a defined energy), and the efficiency of the readout. Following the guidelines presented in the NEMA standards for gamma camera qualifications [NEMA2023], we performed sensitivity measurements, ensuring that rates did not exceed 20’000 counts per second (cps).

### v2

@@ Abstract
Background: The POSiCS camera is a handheld and wireless, small field-of-view gamma camera developed for use in radio-guided surgery (RGS), with sentinel lymph node biopsy (SLNB) as its benchmark application. This compact and lightweight detector (about ) maps tissues labeled with a radiotracer to highlight target lesions. By enabling intraoperative visualization close to the surgical field, POSiCS aims to reduce invasiveness and operative time, improve localization accuracy, and minimize post-operative complications.
Methods: The performances of two hardware configurations were tested: high-sensitivity and high-resolution, to meet application-specific needs. Performance characterization employed Flood (flat) sources and point sources. The response to was also investigated, given its growing role in theranostics. Tests assessed key parameters, including extrinsic spatial resolution, energy resolution, and sensitivity, as well as shielding effectiveness and response to different gamma energies.
Results: In direct contact, the spatial resolution was in high-sensitivity mode and in high-resolution mode. System sensitivity at source-detector distance was and , respectively. Both modes showed an energy resolution of about 20% at 140 keV, though the high-resolution collimator exhibited increased scatter due to its larger tungsten content.
Conclusion: The POSiCS camera combines compactness, ease of use, and reliable performance, making it suitable for RGS. Its interchangeable configurations allow balancing between sensitivity and resolution depending on surgical needs. The results demonstrate millimeter-scale resolution and adequate sensitivity for and imaging, confirming its suitability for intraoperative use. These outcomes support the camera’s potential for broader clinical deployment and further development toward advanced intraoperative nuclear imaging.
@@
Intraoperative Gamma Camera; Radio-guided Surgery; Sentinel Lymph Node Biopsy; Silicon PhotomultipliersContents
1 Introduction
@@
Radio-guided surgery (RGS), including intraoperative techniques, such as sentinel lymph node biopsy (SLNB) and radioguided occult lesion localization (ROLL), is a widely adopted approach, particularly in the surgical management of breast cancer and cutaneous melanoma. [RGS-breast-cancer, Mariani811]. Despite the availability of well-established alternatives to RGS (such as Fluorescence-Guided Surgery (FGS) or Ultrasound-Guided Surgery), the preoperative administration of radio-labeled molecules remains one of the most widely used techniques for intraoperative localization of oncological lesions. This is due to the large mean free path of gamma rays in living tissues, which allows for the direct detection and localization of activity hot-spots from outside the patient’s body [RGS-breast-cancer, Mariani811, Arabi_2024].
@@
The state-of-the-art practice in SLNB and ROLL is based on radiation-counting probes to localize lesions or sentinel lymph nodes (SLNs) [Tsuchimochi_2013, Farnworth2023]. Although this approach has been proven reliable by clinically comparing preoperative lymphoscintigraphies with the number of SLNs localized through a probe-based search, probes still inherently bear a number of drawbacks. To address this shortcoming, an increasing number of 2D Intraoperative Gamma Cameras (IGCs), also known as Small Field-of-View (SFOV) gamma cameras, have been developed in the last two decades [Stoffels2012, Massari2018]. This trend has been thoroughly studied in recent years, thanks to the comprehensive reviews in the field of IGCs, such as [Tsuchimochi_2013] and a more recent follow-up [Farnworth2023]. The main advantages of these cameras are their imaging capabilities, which allow surgeons to localize hot spots near the ”blast zone” (radiotracer injection site) and to provide a method to verify post-excision images to ensure that the target lymph node or lesion was entirely resected [Hellingman_2016, Paredes2008]. Additionally, imaging is intuitive and does not require the specific and lengthy training needed to become proficient with probes. It has also been shown that 2D imaging of oncological lesions can reduce surgery duration, as it facilitates the localization of target tissues more quickly [Hellingman_2016, soluri2006].
POSiCS (Position-sensitive SiPM Compact and Scalable gamma-Camera) is a handheld, compact, and wireless gamma camera designed for real-time intraoperative guidance. The device is based on a pixelated LYSO:Ce scintillator and a 33 array of position-sensitive SiPMs (LG-SiPMs). This module is equipped with two interchangeable collimators: a short septa, low-energy, high-sensitivity (LEHS) collimator and a longer, low-energy, high-resolution (LEHR) collimator, enabling dual operating modes that can be selected intraoperatively to adapt to different clinical needs. Designed to effectively detect photons at , corresponding to the emission energy of [KleinJan2013], the standard radionuclide used in RGS, the POSiCS camera exhibits enhanced performance relative to existing systems.
@@
In this work, we present the results of the first performance studies on the prototype POSiCS camera module. In Section 2, we present the experimental methods adopted to measure the main performance parameters of the camera with both collimators. The tests were inspired from the 2023 NEMA standards for gamma camera qualification [NEMA2023]. However, as these standards are primarily designed for large clinical scanners, only a subset of the proposed measurements and procedures applies to our device, given its specific design and use case. In Section 3, the results of the tests are displayed and commented. Conclusively, in Section 4, we provide an overall discussion of the obtained results, while in Section 5, we offer a brief outlook and prospects.
2 Materials and Methods
2.1 The POSiCS camera
@@
The current POSiCS camera module has a total length of 147 mm (155 mm with the LEHR collimator mounted), with a Field-of-View (FOV) of . From this, we define a Useful FOV (UFOV) of . The module total weight is 313 g (with the LEHS collimator) and 381 g (with the LEHR collimator). Further technical details about the camera are provided elsewhere [Acerbi2025]. A picture of the camera’s prototype wireless module is shown in Figure 1.
2.1.1 Collimator geometry
@@
The POSiCS camera is equipped with two interchangeable parallel-hole collimators: one targeting high spatial resolution (LEHR) for surgical procedures that require precise identification of target lesion margins, and the other designed to achieve high sensitivity (LEHS) for faster scanning, even at low injected activities. The different parameters that qualify both objects were chosen through theoretical calculations and simulations using GATE [GATE2004]. Due to its high atomic number (implying a short attenuation length) and good resistance to shocks [Farnworth2023], tungsten was chosen. All collimator parameters were chosen to ensure a septal penetration probability of less than [collimator-design, cherry_ch14]. The collimator was designed with external shielding walls that extend beyond the collimator holes and support the scintillator, reducing the amount of gamma rays entering from the side of the camera.
2.1.2 LYSO scintillator approach
@@
The radiation-sensitive component of the POSiCS camera is a pixelated LYSO:Ce scintillator. This Lutetium-based crystal is typically used in Positron Emission Tomography (PET) scanners due to its fast decay time (), high light yield (), and high density [Kimble_2002]. The camera was therefore built using a very thin crystal (a few millimeters thick), keeping a high photoelectric absorption probability at . The scintillator is pixelated in a 3030 matrix. Each pixel is coated with a BaSO4 reflector to increase light collection on the light detector.
@@ 2.1.3 LG-SiPMs photodetector
In the current context of light detectors for IGCs, POSiCS proposes an innovative approach by moving towards PS-SiPMs, leveraging the new Linearly Graded SiPMs technology (LG-SiPMs) designed and produced by FBK [Acerbi2024, FBK-LG-SiPM, Acerbi2024_ps]. Through an array of 3x3 LG-SiPMs with peak sensitivity at , an active area of about per sensor, and microcell pitch of , it is possible to reconstruct the position of each scintillation event, through a weighted charge (and amplitude) distribution towards the output channels. Each Single Photon Avalanche Diode (SPAD) composing the SiPM is connected to a resistive network, which performs a light spot CoG reconstruction, encoded in the relative amount of charge output by each readout channel [Acerbi2025, Acerbi2024]. This Anger CoG reconstruction is typically used between SiPMs or PMTs in standard gamma cameras, but this approach is implemented at the single microcell level within the POSiCS camera. The 33 array configuration, with the ”smart channel” approach (see [Acerbi2024]), allows CoG reconstruction with only 8 readout channels: 4 encoding for the x coordinate () and 4 reconstructing the y coordinate (). More precisely, position reconstruction is performed as follows:
This weighted position calculation is performed directly within the camera digital electronics, which in turn sends the reconstructed positions through a wireless connection. This photon detector demonstrated exceptional intrinsic precision, achieving a resolution of a few hundred microns [Acerbi2024, Gola2013ANA, raiola2025spatialresolution].
@@
2.1.4 Approach to evemt reconstruction
After a scintillation event in the scintillator, the eight SiPM signals are amplified by transimpedance amplifiers (TIAs). Their summed charge output is sent to a comparator, which enables charge integration only when a user-defined threshold is crossed. The resulting per-channel charge is integrated and digitized by an eight-channel ADC. The microcontroller (C) then computes the event energy (sum of all channels output charge), timestamp , and the - and -coordinates using Equation 1, and also provides slow-control (SC) data such as bias voltage, current, temperature, and battery status (see Figure 2).
Each event is stored as and transmitted to an external computer via an ESP32 module in packets of 4000 events or after a user-defined time window has elapsed. The data is retrieved via HTTP GET requests. The camera settings are configured through the same ESP32 using HTTP POST commands to the DACs. For further technical details about the POSiCS camera electronics, see [Acerbi2025].
Reconstructed images are corrected using a three-step calibration and correction procedure. First, once the incident position of a gamma ray is determined (x,y CoG coordinates), the event is assigned to the corresponding camera pixel (scintillator crystal) associated with the scintillator crystal encompassing the interaction position. This spatial mapping is established through a flood-field calibration scan, which serves as the primary calibration step before operating the camera.
Second, a pixel-by-pixel energy calibration is performed. In this step, a scaling coefficient is applied to convert the SiPMs charge response into a physical energy. Due to non-uniform light collection across certain regions of the FOV, as well as slight variations in the gain and response characteristics of the SiPMs forming the optical readout array, each pixel is assigned an individual conversion coefficient to convert the digitized charge to a physical energy in keV. Following this step, a user-defined energy window can be applied in software to selectively retain events consistent with photoelectric interactions, thereby suppressing contributions from Compton-scattered events that occur either within the LYSO crystal or in the surrounding environment.
@@ 2.1.4 Approach to evemt reconstruction
With this acquisition system, it is possible to display images live with a delay of the order of 1 second. The final images are enhanced using a 2D-Gaussian kernel convolution, following the guidelines in the 2023 NEMA standards for the validation of gamma cameras [NEMA2023]. Finally, the smoothed image is interpolated using splines of order 2.
2.2 Performance evaluation
@@
To assess spatial resolution, a point source, defined as having a diameter smaller than that of the scintillator pixel, was necessary. Due to the practical limitations of producing such a source with , a solid source was employed. This isotope emits gamma-rays at 122 keV, serving as a suitable surrogate for the 140.5 keV photons of .
@@
2.2.1 Extrinsic Spatial resolution
@@
The uncertainty on the resolution is the summation in quadrature of the statistical error (obtained through the covariance matrix of the fit) and a systematic uncertainty. Systematics is defined as the variation in resolution due to the different behavior of the camera in other areas of the FOV. The systematic uncertainty is evaluated through a scan of 15 positions on the device’s FOV with the 57Co point source, acquiring 100’000 events at each position. Each reconstructed spot is fitted with a Gaussian profile, from which the FWHM is extracted. The standard deviation of this sample of 15 measured resolution values represents the systematic error of the resolution measurement.
The measurement at the center of the FOV is repeated for different source distances and under two conditions: with and without scattering material between the source and the camera head. The chosen scattering material is Polymethyl methacrylate (PMMA). This approach enables us to estimate the expected resolution in biological tissues, whose density is comparable to that of PMMA. The obtained resolution is also compared with the theoretical spatial resolution [cherry_ch14, AngerCamera, xcom2010]:
2.2.2 Sensitivity
The camera’s sensitivity is defined as the registered trigger rate divided by the activity of the source within the camera’s FOV. This parameter is strongly dependent on the geometry of the collimator, the efficiency of the scintillator crystal (i.e., the probability of absorbing a gamma ray through photoelectric effect at a defined energy), and the dead-time of the readout. Following the guidelines presented in the NEMA standards for gamma camera qualifications [NEMA2023], we performed sensitivity measurements, ensuring that rates did not exceed 20’000 counts per second (cps).